OpenStudy (angel010):

Factor completely x2 − 8x + 16. (x + 4)(x + 4) (x − 4)(x − 4) (x + 4)(x − 4) (x − 2)(x − 8)

OpenStudy (angel010):

@andrewhaze

OpenStudy (angel010):

@AlexandervonHumboldt2

OpenStudy (angel010):

@Directrix @HWBUSTER00

Directrix (directrix):

Name two numbers that multiply to 16 and add to -8.

OpenStudy (angel010):

4*4 and 1*16 and 1*-8

Directrix (directrix):

x2 − 8x + 16 = x^2 -4x -4x +16 = x(x -4) -4( x - 4) = (x - 4) * (x - 4) Factoring by grouping.

OpenStudy (angel010):

thankz

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

@Directrix is right